# UltraWiki: Ultra-fine-grained Entity Set Expansion with Negative Seed Entities

### v1

@@ Abstract.
Entity Set Expansion (ESE) aims to identify new entities belonging to the same semantic class as a given set of seed entities. Traditional methods primarily relied on positive seed entities to represent a target semantic class, which poses challenge for the representation of ultra-fine-grained semantic classes. Ultra-fine-grained semantic classes are defined based on fine-grained semantic classes with more specific attribute constraints. Describing it with positive seed entities alone cause two issues: (i) Ambiguity among ultra-fine-grained semantic classes. (ii) Inability to define “unwanted” semantic. Due to these inherent shortcomings, previous methods struggle to address the ultra-fine-grained ESE (Ultra-ESE). To solve this issue, we first introduce negative seed entities in the inputs, which belong to the same fine-grained semantic class as the positive seed entities but differ in certain attributes. Negative seed entities eliminate the semantic ambiguity by contrast between positive and negative attributes. Meanwhile, it provide a straightforward way to express “unwanted”. To assess model performance in Ultra-ESE and facilitate further research, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE. UltraWiki encompasses 50,973 entities and 394,097 sentences, alongside 236 ultra-fine-grained semantic classes, where each query of them is represented with 3-5 positive and negative seed entities. A retrieval-based framework RetExpan and a generation-based framework GenExpan are proposed to comprehensively assess the efficacy of large language models from two different paradigms in Ultra-ESE. Moreover, we devised three strategies to enhance models’ comprehension of ultra-fine-grained entities semantics: contrastive learning, retrieval augmentation, and chain-of-thought reasoning. Extensive experiments confirm the effectiveness of our proposed strategies and also reveal that there remains a large space for improvement in Ultra-ESE.
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/THUKElab/UltraWiki.
1. Introduction
@@
Entity Set Expansion (ESE) is a critical task aiming to identify new entities belonging to the same semantic class as a given set of seed entities (Zhang et al., 2020; Li et al., 2022a, 2023d, 2023g). ESE plays an important role for a wide range of user-tailored applications (Li et al., 2023b; Yu et al., 2023; Li et al., 2022c; Zhang et al., 2023b), such as recommendation systems (Jacucci et al., 2021; Ma et al., 2022; Ye et al., 2023), knowledge base population (Zhu et al., 2023; Velardi et al., 2013; Li et al., 2023a, f, e) and domain text analysis (Kohita et al., 2020; Liu et al., 2022; Li et al., 2023c, 2022b), thus demanding a high degree of semantic class granularity.
@@
Traditional methods of ESE primarily relied on positive seed entities to represent a target semantic class. However, solely leveraging positive seed entities poses a significant challenge in representing ultra-fine-grained semantic classes. Fine-grained semantic classes are concrete at the concept level, e.g., mobile phone brands, car brands, and the target entities of different semantic classes are typically quite different. In contrast, ultra-fine-grained semantic classes are defined based on concepts with more specific attribute constraints, such as mobile phone brands using Android, Asian mobile phone brands. Notably, ultra-fine-grained semantic classes within the same fine-grained class often demonstrate significant overlap in target entities. As shown in Figure 1, describing the target semantic class merely with positive seed entities cause two issues: (i) Ambiguity among ultra-fine-grained semantic classes. For instance, distinguishing between Mobile phone brands using Android and American Mobile phone brands using only positive seeds like {Motorola, Microsoft Mobile, Google} may lead to ambiguous expansion results. These seeds do not capture the distinct features that delineate each class. (ii) Inability to define “unwanted”. Positive seeds alone cannot represent the “unwanted” semantic. Considering when users wish to find Mobile phone brands not made by Asian manufacturers, it is impractical to define this semantic class by enumerating phone brands from other continents (Europe, North America, Africa, etc.) in the positive seeds. Since the size of the positive seeds is limited, and it’s also unreasonable to expect users to specify all potential attribute values.
Figure 1. Semantic classes at different levels of granularities, and issues with positive seed entity only.
Previous exploratory works mainly made efforts towards fine-grained expansion. ProbExpan (Li et al., 2022a) refines fine-grained semantic boundaries by mining in-batch hard negative samples. FGExpan (Xiao et al., 2023) guides the expansion process to yield more precise results for the fine-grained semantic class with the existing taxonomy. MultiExpan (Li et al., 2023g) leverages the visual information of entities to differentiate between semantic classes. Nonetheless, these methods primarily concentrate on conceptual level distinctions (Although these concepts are somehow hard to distinguish, e.g., US states vs. US cities). They are insufficient for distinguishing ultra-fine-grained semantic classes, due to the inherent shortcoming of representing semantic classes with positive seeds only.
@@
Inspired by exploiting the negative feedback to improve the modeling of user purchase intent in the recommendation system (Zhao et al., 2018; Wu et al., 2020; Xie et al., 2021), we solve ultra-fine-grained ESE (Ultra-ESE) by introducing negative seed entities in the inputs. Specifically, the input negative seed entity belongs to the same fine-grained semantic class as the positive seed entity but differs in certain attributes. Positive and negative seed entities jointly represent and constrain target semantic classes, which addresses the representation granularity challenge head-on: Firstly, the contrast between positive and negative seeds on attributes can express the attribute of user interests, eliminating the semantic class ambiguity caused by positive seeds alone. For instance, for the aforementioned example, when the negative seeds are entities of Mobile phone brands using iOS, it indicates the user’s concern about the operation system, irrespective of the manufacturer. Secondly, negative seeds provide a straightforward way to express “unwanted” semantics.
@@
Regrettably, current ESE dataset (Shen et al., 2017; Pradhan et al., 2013) lacks the concept of negative seed entities, let alone ultra-fine-grained semantic classes. To bridge this gap, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE. Derived from Wikipedia, UltraWiki encompasses 10 fine-grained semantic classes, alongside 50,973 entities and 394,097 sentences. We determined 2-3 attributes for each fine-grained semantic class and annotated each entity with attribute values utilizing a semi-automatic approach. Based on the identified attributes, we further constructed 236 ultra-fine-grained semantic classes, in which each semantic class is represented with 3-5 positive and negative seed entities, respectively. In other words, there is an average of 23 (236/10-1) hard negative semantic classes for each ultra-fine-grained semantic class, which poses a significant challenge to the model’s capability for ultra-fine-grained semantic comprehension. Even advancing GPT-4 can’t handle it well.
@@
In experiments, we evaluate the performance of existing methods on UltraWiki. Furthermore, we designed a retrieval-based framework with Encoder-only models (BERT (Devlin et al., 2018)) and a generation-based framework with Decoder-only models (LLaMA (Touvron et al., 2023)) to comprehensively assess the efficacy of large language models (LLMs) from two different paradigms on the Ultra-ESE task. Meanwhile, we proposed three strategies to enhance models’ ability to comprehend ultra-fine-grained semantics of entities: contrastive learning, retrieval augmentation, and chain-of-thought reasoning.
In summary, the main contributions include:
@@
  • We proposed the more challenging ultra-fine-grained ESE task, and introduce negative seed entities to represent ultra-fine-grained semantic classes more precisely for the first tiem.
@@
  • We designed both retrieval-based and generation-based frameworks to assess the efficacy of BERT-like and GPT-like LLMs on the UltraWiki dataset. Furthermore, three strategies are proposed for refining the semantic comprehension of ultra-fine-grained entities.
@@
  • Extensive experiments confirm the effectiveness of our proposed strategies and also reveal significant potential for enhancing ultra-fine-grained semantic comprehension of entities by LLMs.
2. Related Work
2.1. Methods of ESE
@@
Recently, increasing emphasis has been placed on the expansion of fine-grained (e.g., US city) rather than coarse-grained (e.g., Location) semantic classes. ProbExpan (Li et al., 2022a) employs heuristic methods to mine hard negative entities of target semantic classes, thereby refining the semantic boundaries of fine-grained classes through contrastive learning (Chen et al., 2020; Gao et al., 2021). FGExpan (Xiao et al., 2023) leverages the existing taxonomy to direct BERT in reasoning about more fine-grained names of semantic classes. Besides, MultiExpan (Li et al., 2023g) integrates multimodal pre-trained models to encode visual information (images) associated with entities, which benefits the differentiation of various fine-grained semantic classes. Nevertheless, these methods do not essentially tackle the issue that solely using positive seed entities fails to sufficiently represent ultra-fine-grained semantic classes. In contrast, our research introduces the novel concept of incorporating negative seed entities alongside positive ones to describe an ultra-fine-grained semantic class.
@@
It is worth mentioning that there has been some work to incorporate negative entities (Curran et al., 2007; Jindal and Roth, 2011; Shi et al., 2014; Li et al., 2022a), though the usage of negative entities in these methods is relatively naive. We point out that the role of the negative entities used in previous work is fundamentally different from ours. Negative entities in previous work are purely used to help determine the boundary of the target set described by positive seeds. In contrast, negative entities in our model are used to describe the target ultra-fine-grained classes that cannot be characterized by positive seeds alone.
2.2. Data Resources of ESE
@@
Queries in the existing ESE dataset merely consist of positive seed entities. The approach of representing ultra-fine-grained semantic classes solely via positive seed entities inherently faces two major issues: (i) Ambiguity among ultra-fine-grained semantic classes. (ii) Inability to define “unwanted” semantics. Consequently, existing datasets lack ultra-fine-grained semantic classes. For instance, Wiki (Shen et al., 2017) and APR (Shen et al., 2017), derived from the Wikipedia and Reuters corpora respectively, encompass merely 3 and 8 fine-grained semantic classes. Some named entity recognition datasets are directly used as ESE evaluation benchmarks, they contain semantic classes with coarser granularity. For example, the OntoNotes (Pradhan et al., 2013) and CoNLL (Sang and De Meulder, 2003) datasets include semantic classes like organization and location. In contrast, UltraWiki encompasses 261 ultra-fine-grained semantic classes, which surpasses existing ESE datasets in term of both granularity and quantity of semantic classes.
Figure 2. llustration of the UltraWiki dataset construction process.
Figure 3. Distribution of positive and negative target entities in two different cases.
3. Task Formulation
@@
Our study focuses on the ultra-fine-grained ESE task, whose input incorporates three components: query $S$, candidate entity vocabulary $V$, and corpus $D$. The query $S=\{S^{pos}\cup S^{neg}\}$ is a composite set comprising positive seed entities $S^{pos}=\{e^{pos}_{1},\ldots,e^{pos}_{k}\}$ and negative seed entities $S^{neg}=\{e^{neg}_{1},\ldots,e^{neg}_{k}\}$, which jointly represent a target ultra-fine-grained semantic class. Both $S^{pos}$ and $S^{neg}$ belong to the same fine-grained semantic class. The entities in $S^{pos}$ and $S^{neg}$ share same values for certain attributes $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$, respectively. Notably, as shown in Figure3, $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$ are not limited in quantity and may overlap or differ. The vocabulary $V$ enumerates the candidate entities. Additionally, the corpus $D$ supplies contextual information $\{e_{i},s^{i}_{1},\ldots,s^{i}_{n}\}$ for each entity $e_{i}$, in which $s^{i}_{n}$ is a sentence containing entity $e_{i}$.
@@
The ultimate objective of ultra-fine-grained ESE is to identify positive target entities from the vocabulary $V$ that belong to the same fine-grained semantic class, and share the same attribute values with $S^{pos}$ while distinct from $S^{neg}$. In an ideal entity feature space, entities that share more same attribute values shall be positioned closer.
4. UltraWiki Dataset
@@
Considering the lack of negative seed entities and ultra-fine-grained semantic classes in existing ESE datasets, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE. This section outlines the construction process of UltraWiki. Semantic granularity, annotation consistency, and candidate entities’ difficulty are fully considered to guarantee the quality of UltraWiki.
4.1. Dataset Construction
@@
Two methods exist for constructing UltraWiki. A more direct method is to first collect a large number of entities and corresponding context sentences. Then, for each fine-grained semantic class, the human annotator traverses the entire entity vocabulary to identify the corresponding entity and annotates the attribute values, thereby constructing the ultra-fine-grained semantic class. However, despite the availability of numerous entity-centric datasets, this method is labor-intensive and prone to omissions of entities. Consequently, we chose a more practical strategy. We begin by identifying fine-grained semantic classes along with their corresponding entities, then collecting corresponding sentences and attribute information for each entity in turn. Figure 2 illustrates the entire process.
@@
Step 1. Semantic Classes and Entities Collection. Wikipedia maintains an extensive catalog of entities associated with fine-grained semantic classes. We carefully chose 10 semantic classes with certain principles (discussed in Section 3.2) and then crawled the corresponding entities. Additionally, a substantial number of entities sampled from Wikipedia pages were incorporated into the candidate entity vocabulary as negative entities.
@@
Step 2. Entity-Labeled Sentences Collection. A significant volume of text is also crawled from Wikipedia pages, where entities are uniquely identified by hyperlinks. Since the entities obtained in Step 1 also contain hyperlinks, we can readily establish associations between entities and the sentences containing them, thereby providing abundant contextual information about the entities.
@@
Step 3. Entity Attribute Annotation. For each fine-grained semantic class $C$, we manually select $k$ independent attributes $\mathcal{A}=\{a_{1},a_{2},\cdots,a_{k}\}$, which will be used to generate ultra-fine-grained semantic class in next step. Attributes of each class are shown in appendix. Subjective attributes are avoided like good or bad. We initially employed a script to automatically query Wikidata for the corresponding attribute value. For the remaining attributes that could not be automatically annotated, we employed professional annotators to manually label them, ensuring each entity was reviewed three times for accuracy and consistency.
@@
Step 4. Negative-aware Semantic Classes Generation. We devised an algorithm for automated generation of ultra-fine-grained semantic classes based on positive and negative attributes. Specifically, for attributes $\mathcal{A}$ of each fine-grained semantic class $C$, we sample $n^{pos}$ and $n^{neg}$ attributes as positive and negative attributes (denoted as $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$). Subsequently, we pick a value for all the attributes in $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$. This process yields a list of positive values $\mathcal{V}^{pos}$ and a list of negative values $\mathcal{V}^{neg}$, which jointly constrain the attributes of the target semantic class to achieve ultra-fine-grained control of the semantic class. When $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$ are the same, the role of $\mathcal{A}^{neg}$ is to emphasize attributes concerned by user and mitigate ambiguity. Conversely, when they differ, $\mathcal{A}^{neg}$ serves to convey unwanted semantics. We denote the set of entities satisfying $\mathcal{V}^{pos}$ and $\mathcal{V}^{neg}$ as positive target entity $\mathcal{P}$ and negative target entity $\mathcal{N}$ respectively, ensuring that both $|\mathcal{P}|$ and $|\mathcal{N}|$ exceed the minimum entity requirement $n_{thred}=6$.
Figure 4. Heat map visualization of semantic class similarity.
4.2. Dataset Analysis

### v3

@@ Abstract
Entity Set Expansion (ESE) aims to identify new entities belonging to the same semantic class as the given set of seed entities. Traditional methods solely relied on positive seed entities to represent the target fine-grained semantic class, rendering them tough to represent ultra-fine-grained semantic classes. Specifically, merely relying on positive seed entities leads to two inherent shortcomings: (i) Ambiguity among ultra-fine-grained semantic classes. (ii) Inability to define “unwanted” semantics. Hence, previous ESE methods struggle to address the ultra-fine-grained ESE (Ultra-ESE) task. To solve this issue, we first introduce negative seed entities in the inputs, which jointly describe the ultra-fine-grained semantic class with positive seed entities. Negative seed entities eliminate the semantic ambiguity by providing a contrast between positive and negative attributes. Meanwhile, it provides a straightforward way to express “unwanted”. To assess model performance in Ultra-ESE and facilitate further research, we also constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE. UltraWiki encompasses 50,973 entities and 394,097 sentences, alongside 236 ultra-fine-grained semantic classes, where each class is represented with 3-5 positive and negative seed entities. Moreover, a retrieval-based framework RetExpan and a generation-based framework GenExpan are proposed to provide powerful baselines for Ultra-ESE. Additionally, we devised two strategies to enhance models’ comprehension of ultra-fine-grained entities’ semantics: contrastive learning and chain-of-thought reasoning. Extensive experiments confirm the effectiveness of our proposed strategies and also reveal that there remains a large space for improvement in Ultra-ESE. All the codes, dataset, and supplementary notes are available at https://github.com/THUKElab/UltraWiki.
Index Terms:
Entity Set Expansion, Knowledge Mining, Language Model, Ultra-fine-grained Semantic Understanding
I Introduction
Figure 1: The figure illustrates how relying only on positive seeds can lead to ambiguity between similar classes (e.g., Mobile phone brands using Android vs. American Mobile phone brands) and the inability to effectively define negative constraints (e.g., Mobile phone brands not made by Asian manufacturers). The introduction of negative seed entities in the input solves these issues.
@@ I Introduction
Entity Set Expansion (ESE) is a critical task aiming to expand new entities belonging to the same semantic class as the given set of seed entities [1, 2, 3, 4]. For example, given the input seed entity set {Motorola, Xiaomi, Nokia}, an ESE model will output more new entities (e.g., “Huawei”, “Samsung”, …) that all belong to the same Mobile Phone brands class as the seed entities. ESE plays an important role for a wide range of user-tailored applications [5, 6, 7, 8, 9], thus demanding a high degree of semantic class granularity. For instance, in the context of recommendation systems [10, 11], a more nuanced ESE method can benefit more precise product recommendations.
Conventionally, existing ESE methods focus on the expansion of “wanted” fine-grained semantic classes merely based on positive seed entities, such as Car Brand and the above mentioned Phone Brand. However, current ESE methods struggle to expand ultra-fine-grained semantic classes [12, 13, 14], which involve more specific attribute constraints (e.g., mobile phone brands using Android, Asian mobile phone brands). Since ultra-fine-grained semantic classes from the same concept often exhibit substantial overlap in target entities, rendering them tough to be represented only with a handful of positive seed entities.
@@
Specifically, as shown in Figure 1, merely relying on positive seed entities to represent ultra-fine-grained semantic classes causes two issues: (i) Ambiguity among ultra-fine-grained semantic classes. For instance, with the positive seed entities {Motorola, Microsoft Mobile, Google}, it is confusing whether the user intends to extend Mobile phone brands using Android or American Mobile phone brands. Since the positive seed entities alone are insufficient to delineate the distinct class features. (ii) Inability to define “unwanted”. Positive seed entities alone fail to represent the “unwanted” semantic. Considering when users wish to find more Mobile phone brands not made by Asian manufacturers entities, it is impractical to define this semantic class by enumerating brands from other continents in positive seed entities. As the size of the positive seed entities is limited, it’s also unreasonable to expect users to specify all potentially desired attribute values.
@@
Due to the inherent shortcomings of the task inputs, prior ESE methods failed to conduct ultra-fine-grained expansion. To address this issue, inspired by exploiting negative feedback to model user intent in recommendation system [15, 16, 17], we propose enhancing ESE by incorporating negative seed entities, aiming at ultra-fine-grained ESE (Ultra-ESE). The input negative seed entities belong to the same fine-grained semantic class as the positive seed entities but differs in certain attributes. They jointly represent target semantic class, addressing the semantic granularity issue head-on: First, the contrast between positive and negative seed entities in terms of attributes highlights the user’s specific interests and eliminates the ambiguity arising from positive seed entities alone. For the aforementioned example, if negative seed entities are about Mobile phone brands not using Android, it indicates that the user is focused on operating system rather than manufacturer. Second, negative seed entities constrain the expansion space and can readily express “unwanted” semantics.
@@
Regrettably, previous ESE dataset [18, 19] lacked the concept of negative seed entities, let alone ultra-fine-grained semantic classes. To bridge this gap, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE. Derived from Wikipedia, UltraWiki encompasses 10 fine-grained semantic classes, 50,973 entities, and 394,097 sentences. During the curation, we annotated 2-3 attributes for each fine-grained semantic class to further construct ultra-fine-grained semantic classes. Based on the permutations of attributes, 235 ultra-fine-grained semantic classes are constructed, with each semantic class represented with 3-5 positive and negative seed entities. On average, each ultra-fine-grained semantic class contains 23 (236/10-1) hard negative semantic classes, which belong to the same fine-grained semantic class and may exhibit substantial overlap in target entities. Experiments proved that even advancing GPT-4 can not handle it well.
@@
In experiments, we evaluated existing ESE methods on UltraWiki. Furthermore, to comprehensively assess the efficacy of large language models (LLMs) with two different paradigms on the Ultra-ESE task, we designed a retrieval-based framework, RetExpan, with encoder-only LLM BERT [20] and a generation-based framework, GenExpan, with decoder-only LLM LLaMA [21]. Meanwhile, we proposed three strategies to enhance models’ ability to comprehend ultra-fine-grained semantics of entities: contrastive learning, retrieval augmentation, and chain-of-thought reasoning.
In summary, the main contributions are listed as follows:
@@
  • We proposed the more challenging ultra-fine-grained ESE task, and incorporated negative seed entities to represent ultra-fine-grained semantic classes more precisely for the first time.
@@
  • We designed both retrieval-based and generation-based frameworks to assess the efficacy of BERT-like and GPT-like LLMs on the UltraWiki dataset. Furthermore, two strategies were proposed for refining the semantic comprehension of ultra-fine-grained entities.
@@
  • Extensive experiments confirmed the effectiveness of our proposed strategies and also reveal significant potential for enhancing ultra-fine-grained semantic comprehension of entities by LLMs.
II Related Work
II-A Methods of ESE
@@
Recently, increasing emphasis has been placed on the expansion of fine-grained (e.g., US city) rather than coarse-grained (e.g., Location) semantic classes. ProbExpan [3] employs heuristic methods to mine hard negative entities of target semantic classes, thereby refining the semantic boundaries of fine-grained classes through contrastive learning [22, 23]. FGExpan [24] leverages the existing taxonomy to direct BERT in reasoning about more fine-grained names of semantic classes. Besides, MultiExpan [4] integrates multimodal pre-trained models to encode visual information (images) associated with entities, which benefits the differentiation of various fine-grained semantic classes. Nevertheless, these methods do not essentially tackle the issue that solely using positive seed entities fails to sufficiently represent ultra-fine-grained semantic classes. In contrast, our research introduces the novel concept of incorporating negative seed entities alongside positive ones to describe an ultra-fine-grained semantic class.
@@
It is worth mentioning that there has been some work to incorporate negative entities [25, 26, 27, 3], though the usage of negative entities in these methods is relatively naive. We point out that the role of the negative entities used in previous work is fundamentally different from ours. Negative entities in previous work are purely used to help determine the boundary of the target set described by positive seed entities. In contrast, negative entities in our model are used to describe the target ultra-fine-grained classes that cannot be characterized by positive seed entities alone.
II-B Data Resources of ESE
@@
Queries in the existing ESE dataset merely consist of positive seed entities. The approach of representing ultra-fine-grained semantic classes solely via positive seed entities inherently faces two major issues: (i) Ambiguity among ultra-fine-grained semantic classes. (ii) Inability to define “unwanted” semantics. Consequently, existing datasets lack ultra-fine-grained semantic classes. For instance, Wiki [18] and APR [18], derived from the Wikipedia and Reuters corpora respectively, encompass merely 3 and 8 fine-grained semantic classes. Some named entity recognition datasets are directly used as ESE evaluation benchmarks, they contain semantic classes with coarser granularity. For example, the OntoNotes [19] and CoNLL [28] datasets include semantic classes like organization and location. In contrast, UltraWiki encompasses 261 ultra-fine-grained semantic classes, which surpasses existing ESE datasets in terms of both granularity and quantity of semantic classes.
III Task Formulation
@@
We focus on the ultra-fine-grained ESE task, whose input consists of three components: query $S$, candidate entities $V$, and corpus $D$. The query $S=\{S^{pos}\cup S^{neg}\}$ is a composite set comprising positive seed entities $S^{pos}=\{e^{pos}_{1},\ldots,e^{pos}_{k}\}$ and negative seed entities $S^{neg}=\{e^{neg}_{1},\ldots,e^{neg}_{k}\}$, which belong to the same fine-grained semantic class $c$. The entities in $S^{pos}$ share the same values for positive attribute set $\mathcal{A}^{pos}$:
$e.\operatorname{get}(a_{i})\equiv k_{i},\forall e\in S^{pos},\forall a_{i}\in\mathcal{A}^{pos}=\left\{\left(a_{i},k_{i}\right)\right\}_{i=1}^{|\mathcal{A}^{pos}|}$ (1)
where $a_{i}$ and $k_{i}$ are attribute and value, respectively. We define positive target entities $\mathcal{P}$ as entities that share the same attribute values with positive seed entities in $\mathcal{A}^{pos}$. Likewise, negative target entities in $\mathcal{N}$ share the same attribute values with negative seed entities in $\mathcal{A}^{neg}$.
Overall, the target ultra-fine-grained semantic class means class $c$ that holds the same values with positive seed entities in $\mathcal{A}^{pos}$, but different values with negative seed entities in $\mathcal{A}^{neg}$. Examples can be found in Figure 2.
@@
Hence, the ultimate objective of ultra-fine-grained ESE is to identify entities from candidate entities $V$ that belong to the $\mathcal{P}$, while being distinct from $\mathcal{N}$ (e.g., $\mathcal{P}-\mathcal{N}$). In an ideal feature space, entities that share more of the same attribute values should be positioned closer.
Figure 2: Left target semantic class: mobile phone brand with Android OS, not produced by Asian manufacturer. Right target semantic class: mobile phone brand with Android OS. The target entity set is $\mathcal{P}-\mathcal{N}$.
IV UltraWiki Dataset
Figure 3: Illustration of the UltraWiki dataset construction process. The intermediate data obtained at each step is bolded in the figure.
@@ IV UltraWiki Dataset
Considering the lack of ultra-fine-grained semantic classes and negative seed entities in existing ESE datasets, we constructed UltraWiki, the first large-scale dataset tailored for Ultra-ESE. This section outlines the construction process of UltraWiki. Semantic granularity, annotation consistency, and candidate entities’ difficulty are fully considered to guarantee the quality of UltraWiki.
IV-A Dataset Construction
@@
There are two strategies for constructing UltraWiki. A more direct strategy is to first collect a large number of entities and corresponding context sentences. Then, for each fine-grained semantic class, the human annotator traverses the entire entity vocabulary to identify the corresponding entity and annotates the attribute values, thereby constructing the ultra-fine-grained semantic class. However, despite the availability of numerous entity-centric datasets, this method is labor-intensive and prone to omit entities. Consequently, we chose a more practical strategy. We begin by identifying fine-grained semantic classes along with their corresponding entities, then collecting corresponding sentences and attribute information for each entity in turn. Figure 3 illustrates the entire process.
@@
Step 1. Semantic Classes and Entities Collection. Wikipedia maintains an extensive catalog of entities associated with fine-grained semantic classes. We identified 10 fine-grained semantic classes and then crawled the corresponding entities. Additionally, a substantial number of entities sampled from Wikipedia pages were incorporated into the candidate entity vocabulary as negative entities.
@@
Step 2. Entity-Labeled Sentences Collection. A significant volume of text is also crawled from Wikipedia pages, where entities are uniquely identified by hyperlinks. Since the entities obtained in Step 1 also contain hyperlinks, we can readily align entities with sentences containing them, thereby providing abundant contextual information for entities.
@@
Step 3. Entity Attribute Annotation. For each fine-grained semantic class $C$, we manually select $k$ independent attributes $\mathcal{A}=\{a_{1},a_{2},\cdots,a_{k}\}$, which will be used to generate ultra-fine-grained semantic class in next step. Attributes of each class are shown in supplementary notes. Subjective attributes are avoided like good or bad. We first query Wikidata API for the corresponding attribute value with Python script. For the remaining attributes that could not be automatically annotated, we employed human annotators to manually label them, ensuring each entity was reviewed three times for accuracy and consistency.
@@
Step 4. Negative-aware Semantic Classes Generation. We devised an algorithm for automated generation of ultra-fine-grained semantic classes based on annotated attributes. Specifically, for attributes $\mathcal{A}$ of each fine-grained semantic class $C$, we sample $m$ and $n$ attributes as positive and negative attribute set (i.e., $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$). Subsequently, we pick a value for all the attributes in $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$. This process yields a tuple of positive values $\mathcal{V}^{pos}$ and a tuple of negative values $\mathcal{V}^{neg}$, which jointly constrain the attributes of the target semantic class to achieve ultra-fine-grained control of the semantic class.
When $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$ are the same, the role of $\mathcal{A}^{neg}$ is to emphasize attributes concerned by user and mitigate ambiguity. Conversely, when they differ, $\mathcal{A}^{neg}$ serves to convey unwanted semantics. As described in the Section III, positive and negative target entity sets are denoted as $\mathcal{P}$ and $\mathcal{N}$. Both $|\mathcal{P}|$ and $|\mathcal{N}|$ are ensured to exceed the minimum entity requirement $n_{thred}=6$.
Figure 4: Semantic similarity heatmap of ultra-fine-grained semantic classes. Rows/columns: averaged embeddings of ground-truth positive entities per class; cell colors: pairwise cosine similarities. Classes were proportionally sampled to 80 classes based on total size for clarity.
IV-B Dataset Analysis